AIT has a staff of 4 cleaners, each working 8 hours per day (for a payroll...

  1. AIT has a staff of 4 cleaners, each working 8 hours per day (for a payroll cost of GH¢ 640/ day) and overhead expenses of GH¢ 400 / day. The cleaners mop and clean 8 floors each day. The company recently purchased a new machine system that will mop and cleans 14 floors per day. Although the cleaners, their works hours, and pay will be same, the overheads expenses are now GH¢ 800 per day. Compute the following;

  1. labour productivity with the old system.
  2. labour productivity with the new system:
  3. multifactor productivity with the old system:
  4. multifactor productivity with the new system:
Answer #1

Answer i :

labor productivity with the old system =8/4

=2 floors/cleaner

------------------------------------

Answer ii :

labour productivity with the new system =14/4

=3.5 floors/cleaner

--------------------------------------

Answer iii :

multifactor productivity with the old system=8/(4*640+400)

=0.0027

-----------------------------------------

Answer iv :

multifactor productivity with the new system: =14/(4*640+800)

=0.0042